# Large-scale Generative AI Models Lack Visual Number Sense

### v1

@@ Discussion
Responses given by VQA models were automatically parsed. If present, number words were converted to numerical values using the word2number Python library. The response was discarded if it contained multiple numbers or if it contained vague quantification terms (e.g., “a few”, “a bunch of”). We checked that at least 20 correct trials were recorded for each number/category. All models achieved the minimum number of tests required without the need to further prompt them (total number of responses discarded for ViLT: 0; BLIP-2: 6).
Images produced by the image generation models were discarded when more than 50% of at least one target object fell outside the margin, or the model generated only objects of non-target categories, or the image contained more than 20 objects. Target objects were not counted when they appeared at the border of the picture and only a very small portion (no more than 20%) of the object was visible, or they were not well-formed (e.g., the shape was clearly distorted) preventing unambiguous identification, or they were occluded by other objects, preventing unambiguous identification. Target objects were considered valid and thus counted when they appeared behind other objects but could still be reliably identified, or they were not entirely visible (e.g., more than 50% of the object mass fell outside the margin or behind other objects) but a salient feature of the object was clearly visible (e.g., the face of a person or the stalk of an apple), or they were represented in an unconventional way (e.g., an apple cut in half).
@@
For DALL-E, the minimum number of trials was not achieved for the dots category with n = 10 (15), since the model in these cases most frequently generated more than 30 items. For Stable Diffusion, the minimum number was not achieved for the dots category with n = 1 (15), n = 8 (14), n = 9 (14) and n = 10 (13). Also in this case, for these prompts the model most frequently generated more than 30 items.
@@ Acknowledgements
We are grateful to OpenAI for granting research access to the DALL-E APIs. This work was partially supported by the Italian Ministry of University and Research (PRIN grant n. C53D23004110006 to M.Z.).

### v2

@@ Discussion
Responses given by VQA models were automatically parsed. If present, number words were converted to numerical values using the word2number Python library. The response was discarded if it contained multiple numbers or if it contained vague quantification terms (e.g., “a few”, “a bunch of”). Interestingly, in a few cases the Gemini model produced unexpected responses with images containing only one item, for example answering that “There are two things in the image: an apple and a white background”. We discarded these answers. We checked that at least 20 correct trials were recorded for each number / category. All models achieved the minimum number of tests required without the need to further prompt them (total number of responses discarded for ViLT: 0; BLIP-2: 6; GPT-4V: 0; Gemini: 20).
Images produced by the image generation models were discarded when the annotators judged them to be too ambiguous to be correctly parsed. The criteria for discarding an image were the following (representative examples are shown in Supplementary Figure SI1):
  • the model generated only objects of non-target categories (Fig. SI1a);
  • the model generated target objects that could in principle be identified as such, but were not well-formed because the shape of the item was significantly distorted and/or other distinguishing features were significantly altered (see examples b and c in Fig. SI1);
  • the model generated one or more objects that resembled the target category but could not be reliably identified as such in isolation (i.e., without providing the entire content of the image as context); this happened, for example, when an object was only partially visible (e.g., less than 10% of the object was included in the frame) and the model did not depict it using representative features, such as the face of a person or the stalk of an apple (see examples d and e in Fig. SI1); vice versa, objects were considered valid and thus counted when they appeared behind other objects or partially outside the image frame, but could nevertheless be reliably identified (see examples f, g and h in Fig. SI1);
  • for the "apple" category, we initially considered valid images containing unconventional representations of the target items (e.g., an apple cut in half); however, DALL-E 3 frequently generated images with entire apples plus several slices of apples and/or additional apples cut in half; to avoid ambiguity we thus opted for discarding those images (see examples i and l in Fig. SI1);
  • for the "people" category, we discarded images mixing realistic people and pictorial content representing people (e.g., an artist painting a portrait, see example m in Fig. SI1), as well as images depicting people in the forefront but also less-visible people in the background (see example n in Fig. SI1);
  • we discarded images containing more than 20 objects (see examples o, p and q in Fig. SI1).
Figure SI1: Examples of images produced by DALL-E 3 that illustrate the criteria we used to discard ambiguous generations. We also report the target number requested in the prompt to highlight the fact that sometimes the discarded image contained the correct amount of items, but often it did not. a) Wrong object category (target category: dots; target number = 6). b) Altered features or distorted object representations for the apples category (target number = 4). c) Altered features or distorted object representations for the butterflies category (target number = 3). d) Unreliable identification of a target object due to the absence of representative features (target number = 4). e) Unreliable identification of an object due to the absence of representative features (target number = 6). f) Reliable identification even though objects are only partially visible (target number = 6; counted 6 objects). g) Reliable identification even though objects are only partially visible (target number = 7; counted 7 objects). h) Reliable identification even though objects are only partially visible (target number = 7; counted 8 objects). i) Apples cut into pieces (target number = 6). l) Apples cut into pieces (target number = 8). m) Pictorial representations of people mixed with real people (target number = 6). n) People in the background (target number = 1). o) More than 20 objects (target number = 8). p) More than 20 objects (target number = 9). q) More than 20 objects (target number = 10).
It is important to clarify that each discarded image was replaced by an additionally generated sample. Moreover, the procedure was blind with respect to the target number of objects. This implies that the performance of a model cannot be penalized by the data cleaning procedure.
@@
For Stable Diffusion, the minimum number was not achieved for the dots category with n = 1 (15), n = 8 (14), n = 9 (14) and n = 10 (13), since the model in these cases most frequently generated more than 30 items. For DALL-E 2, the minimum number of trials was not achieved for the dots category with n = 10 (15). Also in this case, for these prompts the model most frequently generated more than 30 items.
@@ Acknowledgements
We are grateful to OpenAI for granting research access to the GPT-4V and DALL-E APIs. This work was partially supported by the Italian Ministry of University and Research (PRIN grant n. C53D23004110006 to M.Z.).